Hello, I was wondering if anyone could help me with a polynomial function.
It says to find all the real zeros.
f(x)= 2x^4 -2x^2 -40
2(x^4 -x^2 -20) = 0
2(x^2 -5)(x^2 +4) = 0

Answer by stanbon(75887)

Hello, I was wondering if anyone could help me with a polynomial function.
It says to find all the real zeros.
f(x)= 2x^4 -2x^2 -40
2(x^4 -x^2 -20) = 0
Divide both sides by 2 to get:
(x^2 -5)(x^2 +4) = 0
Since the product is zero one of those factors must be zero:
x^2 = 5 or x^2 = -4
x = sqrt5 or x = -sqrt5
---------------
x^2 cannot be -4 if x is a Real Number.
================================
Cheers,
Stan H.